OpenStudy (anonymous):

if cosx=-3/4, find cos2x

OpenStudy (anonymous):

cos2x=2(-3/4)^2-1...from here I a struggling coming up w all different answers

OpenStudy (anonymous):

Thats 2(9/16) - 1 = 18/16 - 1 = 18/16 - 16/16 = 2/16 = 1/8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome.